Two investments earn an annual income of $221. One investment earns an annual simple interest rate of 8.9%, and the other investment earns an annual simple interest rate of 6.6%. The total amount invested is $3000. How much is invested in each account?

Answer by checkley79(3341)

.089X+.066(3,000-X)=221
.089X+198-.066X=221
.023X=221-198
.023X=23
X=23/.023
X=$1,000 INVESTED @ 8.9%.
3,000-1,000=2,000 INVESTED @ 6.6$
PROOF:
.089*1,000+.066*2,000=221
89+132=221
221=221